a 7g bullet moving horizontally at 200m/s strikes and passesthrough a 150 g tin can sitting on a point, just after impact,the can has a horizontal speed of 180m/s.What is the bullet’s speed after hitting the can?
 

Grade:12th pass

1 Answers

Arun
5 years ago
Conservation of momentum. 

mass * velocity of bullet + mass * velocity of can = constant 
Mass of bullet: 7 grams = 0.007 kg 
Mass of can: 150 grams = 0.150 kg 
Initial speed of bullet = 200 m / s 
Initial speed of can = 0 m / s 
Final speed of can = 1.8 m / s 
Final speed of bullet = ?? 
Initial Energy: 
(0.007 kg * 200 m / s) + (0.150 kg * 0 m / s) = 1.4 kg m / sec 
Final Energy: 
0.007 * bullet final speed + 0.150 kg * 1.80 m / s = 1.4 kg m / sec 
0.007 * bullet final speed + 0.27 kg m / s = 1.4 kg m / sec 
bullet final speed = (1.4 - 0.27) / 0.007 = 161.4 m /sec